### v1

@@ Abstract
Continuous-Time Markov Chains Approximate Probabilistic Bisimulation Quasi-Lumpability
@@ 1 Introduction
Continuous-time Markov chains (CTMCs) are a prominent probabilistic model in various application fields, e.g., reliability engineering, systems biology, modeling of chemical reactions, and performance evaluation. CTMCs are state-based models whose transitions yield a discrete probability distribution over states—as in discrete-time Markov chains (DTMCs)—while the state residence times are governed by exponential distributions. Various model-checking approaches for CTMCs exist [7, 11, 22, 63] and are supported by tools such as PRISM [40] and Storm [34]. CTMC model checking is used to analyze e.g., stochastic Petri nets [6], fault trees [8, 61], biological systems [41, 46], and chemical reactions [2, 4].
@@
  • We extend (some of) our results to reward-bounded reachability probabilities in CTMCs with nonnegative rewards, by utilizing a method from [12] that allows to express them as timed reachability probabilities (Section 6).
@@
Uncertain continuous-time Markov chains (UCTMC) [20, 21] are continuous-time Markov chains with time-varying transition rates. At each time $t\geq 0$ and for all states $s,s^{\prime}$, the transition rate to move from $s$ to $s^{\prime}$ must be contained in a fixed interval $[m(s,s^{\prime}),M(s,s^{\prime})]\subseteq\mathbb{R}_{\geq 0}$. Lumpabilities for UCTMCs are equivalences that require related states to have the same “extremal realizations”, defined as the (time-invariant) CTMCs in which all transition rates coincide with $m(s,s^{\prime})$ resp. $M(s,s^{\prime})$. These lumpabilities can be computed efficiently and are characterized both via value functions and satisfaction equivalence of formulas in a variant of the continuous stochastic logic CSL [9, 13]. However, there does not seem to be a direct connection to our notion of $(\varepsilon,\delta)$-bisimulation.
Another related line of research is perturbation theory for Markov chains. Given a chain $\mathcal{M}$ and a (slightly) perturbed $\mathcal{M}^{\prime}$, the goal is to bound the difference of (some) performance measures of the models. See, e.g., [43, 44, 45, 51, 58] or the overview article [52]. The bounds presented in the literature are usually on the total error between the stationary distributions of $\mathcal{M}$ and $\mathcal{M}^{\prime}$. Furthermore, they are oftentimes obtained under the assumption of drift conditions or rely on the ergodicity of the chain. This is in contrast to our work, which focuses on the componentwise difference in transient reachability probabilities up to time $t$.
@@ 2 Preliminaries
A continuous-time Markov chain (CTMC) is a tuple $\mathcal{M}=(S,P,E,s_{init},L)$ such that $\mathcal{D}_{\mathcal{M}}=(S,P,s_{init},L)$ is a DTMC, called the embedded or underlying DTMC of $\mathcal{M}$, and $E\colon S\to\mathbb{R}_{>0}$ is an exit rate function. We use the same notations for CTMC as for DTMC, and we let $\mathcal{M},\mathcal{N}$ range over CTMCs. Note that we do not exclude the possibility of $P(s,s)>0$, i.e., we allow self-loops in CTMCs. The residence time in a state $s$ of $\mathcal{M}$ is negative exponentially distributed with rate $E(s)$, so the probability to take any outgoing transition of $s$ until time $t\geq 0$ (including self-loops) is $1-e^{-E(s)\cdot t}$, and the probability to take a transition from $s$ to some specific state $s^{\prime}$ until $t$ is $P(s,s^{\prime})\cdot(1-e^{-E(s)\cdot t})$. $\mathbf{Q}\in\mathbb{R}^{|S|\times|S|}$ denotes the (infinitesimal) generator of $\mathcal{M}$. The entries of $\mathbf{Q}$ are given as $\mathbf{Q}_{i,j}=P(i,j)\cdot E(i)$ if $i\neq j$ and $\mathbf{Q}_{i,i}=-\sum_{j\neq i}~\mathbf{Q}_{i,j}$. For more information on CTMC and DTMC see, e.g., [14, 36, 39].
@@ 3.2 Definition of $(\varepsilon,\delta)$-Bisimulation
Let $\varepsilon<\frac{1}{2}$ and $\delta>0$. In Figure 2, $\sim_{\varepsilon,\delta}$ is the reflexive and symmetric closure of $\{(s_{0},s_{2}),(s_{0},s_{3}),(s_{1},s_{4}),(s_{2},s_{3})\}$. States $s_{1},s_{2}$ are not $(\varepsilon,\delta)$-bisimilar as they violate the $\delta$-condition: $|\ln(E(s_{1}))-\ln(E(s_{2}))|=2\delta$. Moreover, $s_{0}\nsim_{\varepsilon,\delta}s_{1}$ since although the pair of states satisfies the $\delta$-condition, the $\varepsilon$-condition is violated: $P(s_{0},\{s_{2}\})=\frac{1}{2}>\varepsilon=P(s_{1},{\sim_{\varepsilon,\delta}}(\{s_{2}\}))+\varepsilon.$ Note that the steady state distributions of $(\varepsilon,\delta)$-bisimilar states can differ significantly: starting from $s_{0}$ the probability to reach $g$ (and stay there forever) is $2/3$, from $s_{2}$ it is $1$ and from $s_{3}$ it is 00, even though $s_{0},s_{2}$, and $s_{3}$ are pairwise $(\varepsilon,\delta)$-bisimilar.
@@ 3.3 Fundamental Properties of $\sim_{\varepsilon,\delta}$
By slightly adjusting a procedure proposed in [25] that constructs $\sim_{\varepsilon}$ for DTMC by solving flow networks á la [10, 57], $\sim_{\varepsilon,\delta}$ can be computed in time polynomial in the number of states of $\mathcal{M}$.
@@
Furthermore, $(\varepsilon,\delta)$-bisimilarity in $\mathcal{M}$ implies $\tau$-bisimilarity (for a $\tau\geq 0$ that depends on $\varepsilon,\delta$) in the uniformization $\mathrm{unif}(\mathcal{M},q)$ of $\mathcal{M}$ for $q\geq\max_{s\in S}E(s)$.
@@
For any given $\varepsilon\in(0,1),\delta>0$ and $\tau>0$ there is a CTMC $\mathcal{M}=\mathcal{M}(\varepsilon,\delta,\tau)$ with states $s,s^{\prime}$ such that $\{s,s^{\prime}\}$ is a block of a $\tau$-quasi lumpability on $\mathcal{M}$ but the pair $(s,s^{\prime})$ does not satisfy the $\varepsilon$- nor the $\delta$-condition.
@@
We finish the section by showing how to “split” $(\varepsilon,\delta)$-bisimilarity into $(\varepsilon,0)$- and $(0,\delta)$-bisimilarity. More precisely, given $\mathcal{M}\sim_{\varepsilon,\delta}\mathcal{N}$ we construct CTMCs $\mathcal{M}^{\prime},\mathcal{N}^{\prime}$ with the same graph structure and such that $\mathcal{M}\sim_{\varepsilon,0}\mathcal{M}^{\prime}\sim_{0,\delta}\mathcal{N}^{\prime}\sim\mathcal{N}$. This decomposition makes it possible to treat the two parameters individually and, together with the additivity of $\sim_{\varepsilon,\delta}$, allows us to extend results shown for $(\varepsilon,0)$- and $(0,\delta)$-bisimilar states to $(\varepsilon,\delta)$-bisimilar ones.
@@ 4 Bounding Timed Reachability Probabilities
By using the uniformization $\mathrm{unif}(\mathcal{M},q)$ and applying the bound of Theorem 3.1 [18], which is possible because of the preservation of approximate bisimilarity in the uniformization (see Lemma 2), we derive an upper bound on the absolute difference of timed reachability probabilities of $(\varepsilon,\delta)$-bisimilar states.
@@
A disadvantage of the bounds in Proposition 3 (and Corollary 2) is that they converge to $1$ exponentially fast for $t\to\infty$, while the maximal total difference in timed reachability probabilities is usually much smaller for $(\varepsilon,\delta)$-bisimilar states. Intuitively, the bounds converge to $1$ because the bound of Theorem 3.1 used in their derivation always assumes the maximal possible error. However it does not consider, e.g., that for growing $t$ some of the probability mass already reached the goal state and can thus not contribute to the total error anymore. This disadvantage is particularly observable for $(0,\delta)$-bisimilar states, as in this case the actual error converges to 00 for $t\to\infty$ (see also Figure 7 in Section 5). In Section 5, we derive explicit formulas and better bounds for the absolute difference in timed reachability probabilities of $(0,\delta)$-bisimilar states.
@@ 5 The Effect of Changing Rates Only
Next we show how to obtain, for given $t$ and $\delta$, a maximal $N$ (that depends on $t,\delta$) such that $\mathrm{Diff}_{t}(\mathcal{M})\leq\mathrm{Diff}_{t}(\mathcal{E}_{N})$. Using the explicit form of $\mathrm{Diff}_{t}(\mathcal{E}_{N})$ from Proposition 4, this yields an easy to compute upper bound for $\mathrm{Diff}_{t}(\mathcal{M})$.
@@
Next, we present a way to compute the values of $p_{n}$ based on a spectral decomposition of the transition probability matrix $\mathbf{P}$ of CTMC $\mathcal{M}$. First, we consider the case of diagonalizable $\mathbf{P}$, and afterwards deal with arbitrary $\mathbf{P}$ by using the Jordan canonical form [50]. Moreover, we use the so obtained explicit formulas for $p_{n}$ to obtain upper bounds for $\mathrm{Diff}_{t}(\mathcal{M})$. In contrast to the results so far, these bounds converge to 00 for $t\to\infty$, just like the actual difference does.
The transition probability matrix $\mathbf{P}$ of $\mathcal{M}$ is diagonalizable if there is a diagonal matrix $\mathbf{D}$, with diagonal elements corresponding to the eigenvalues of $\mathbf{P}$ (repeated according to their multiplicities), and a regular matrix $\mathbf{S}$ such that $\mathbf{P}=\mathbf{S}\mathbf{D}\mathbf{S}^{-1}$ [50]. It is well-known that every eigenvalue $\lambda_{i}$ of a stochastic matrix satisfies $|\lambda_{i}|\leq 1$, and that $\lambda_{1}=1$ is an eigenvalue of any such matrix. As we assume that an absorbing state is reached almost surely, $1$ is the only eigenvalue of modulus $1$, and it has a multiplicity $a_{\mathbf{P}}$ of at most $2$ (as there are at most two absorbing states in $\mathcal{M}$). Hence, we can w.l.o.g. assume that, for the $m$ distinct eigenvalues of $\mathbf{P}$ we have $\lambda_{1}=1>|\lambda_{2}|\geq\ldots\geq|\lambda_{m}|$, and that the diagonal of $\mathbf{D}$ is in descending order w.r.t. the absolute values of these eigenvalues. From now on we denote the second largest (absolute value wise) eigenvalue of $\mathbf{P}$ by $\lambda$.
@@
By Proposition 6, $p_{k+1}=\sum_{j=a_{\mathbf{P}}+1}^{n}\mathbf{S}_{1,j}\cdot\mathbf{S}^{-1}_{j,n}\cdot(\lambda_{j}-1)\cdot\lambda_{j}^{k}$, so by the triangle inequality $p_{k+1}~\leq\sum_{j=a_{\mathbf{P}}+1}^{n}|\mathbf{S}_{1,j}\cdot\mathbf{S}^{-1}_{j,n}\cdot(\lambda_{j}-1)|\cdot|\lambda_{j}|^{k}$. Because $|\lambda_{j}|\leq|\lambda|$ for all $j>a_{\mathbf{P}}$ it follows that $p_{k+1}\leq|\lambda|^{k}\cdot\sum_{j=a_{\mathbf{P}}+1}^{n}C=|\lambda|^{j-1}\cdot(n-a_{\mathbf{P}})\cdot C$ for $C=\max_{i=a_{\mathbf{P}}+1,\ldots,n}|\mathbf{S}_{1,i}\cdot\mathbf{S}^{-1}_{i,n}\cdot(\lambda_{i}-1)|$. The claim follows from Theorem 5.1. ∎
@@
Let $\mathcal{M}^{\prime}=e^{0.1}\cdot\mathcal{M}$. In particular, $\mathcal{M}\sim_{0,0.1}\mathcal{M}^{\prime}$. We analyze the bounds from Propositions 5 and 7 w.r.t. $\mathcal{M},\mathcal{M}^{\prime}$ and $f$ by considering three scenarios that are parameterized in $\tau$ and $\mu$ [^2] We omit the bound based on uniformization from Proposition 3 as we have already seen in Example 5 that it does not perform well for $(0,\delta)$-bisimilar models. In the first scenario, $\tau_{1}=\frac{3}{4}$ and $\mu_{1}=\frac{1}{4}$, i.e., the arrival rate of tasks is three times higher than their service rate. In the second scenario, arrival and service rates coincide ($\tau_{2}=\frac{1}{2}=\mu_{2}$), while in the third scenario the service rate is twice the arrival rate, i.e., $\tau_{3}=\frac{1}{3}$ and $\mu_{3}=\frac{2}{3}$. For any $i\in\{1,2,3\}$, the transition probability matrix $\mathbf{P}_{i}$ of the corresponding CTMC is diagonalizable.
@@
We observe that in each scenario the actual error and the bounds of Proposition 7 converge to 00 for $t\to\infty$. The bound of Proposition 5, on the other hand, converges to $1$ for increasing $t$. The convergence rate to 00 of the former bound can, however, be quite slow. In, e.g., the third scenario, the second largest eigenvalue of $\mathbf{P}_{3}$ has modulus $0.9778$, causing a slow convergence of $|\lambda|^{k}$ to 00. On the other hand, in the first scenario $\mathcal{M}_{1}$ enters $f$ quickly, which is reflected in a second largest eigenvalue of $0.7334$ whose powers converge to 00 fast.
While the values of the bounds are, in particular for small $t$, significantly larger than the actual error, it is promising that their shape as a function of $t$ closely resembles the actual difference. This scaling of the bounds by a large factor is caused by the constant $C$ in Proposition 7, which is obtained by using the triangle inequality and a maximum. Refining this constant in future work might lead to bounds following the actual error much more closely.
@@
The decompositions of $\mathbf{P}$ proved in Proposition 6 and Proposition 8 also hold when the Markov chain contains more than $2$ (reachable) absorbing states, i.e, if $a_{\mathbf{P}}>2$. In an absorbing Markov chain, the multiplicity (both geometric and algebraic) of the eigenvalue $1$ is the number of absorbing states, and thus each Jordan block for eigenvalue $1$ has a size of $1$, causing them to get canceled out in the computations of the explicit forms of $p_{N+1}$.
@@
In Theorem 5.2 we have to exclude the case $|\lambda|=0$, i.e., the case that $\mathbf{P}$ only has eigenvalues 00 and $1$. This happens, e.g., if $\mathbf{P}$ (and hence $\mathcal{M}$) is acyclic. In an acyclic CTMC the goal state $g$ is, however, reached after at most as many steps as the length of the longest path in the chain. Therefore, if $\mathcal{M}$ is acyclic, $\mathrm{Diff}_{t}(\mathcal{M})$ can easily be computed exactly, for example by using the identity in Theorem 5.1 and computing the relevant values of $p_{n}=\mathbf{P}^{n+1}-\mathbf{P}^{n}$, or by applying methods like the ACE algorithm of [47] that allow computing the transient distribution functions of acyclic CTMC directly.
@@
6 Bounds for Reward-Bounded Reachability

### v2

@@ Abstract
Continuous-Time Markov Chains Approximate Probabilistic Bisimulation Quasi-Lumpability Time-Bounded Reachability
@@ 1 Introduction
Continuous-time Markov chains (CTMCs) are a prominent probabilistic model in various application fields, e.g., reliability engineering, systems biology, modeling of chemical reactions, and performance evaluation. CTMCs are state-based models whose transitions yield a discrete probability distribution over states—as in discrete-time Markov chains (DTMCs)—while the state residence times are governed by exponential distributions. Various model-checking approaches for CTMCs exist [7, 11, 22, 63] and are supported by tools such as PRISM [40] and Storm [34]. CTMC model checking is used to analyze, e.g., stochastic Petri nets [6], fault trees [8, 61], biological systems [41, 46], and chemical reactions [2, 4].
@@
  • We extend (some of) our results to reward-bounded reachability probabilities in CTMCs with nonnegative rewards, by utilizing a method from [12] that allows us to express them as timed reachability probabilities (Section 6).
@@
Uncertain continuous-time Markov chains (UCTMC) [20, 21] are CTMCs with time-varying transition rates. At each time $t\geq 0$ and for all states $s,s^{\prime}$, the transition rate to move from $s$ to $s^{\prime}$ must be contained in a fixed interval $[m(s,s^{\prime}),M(s,s^{\prime})]\subseteq\mathbb{R}_{\geq 0}$. Lumpabilities for UCTMCs are equivalences that require related states to have the same “extremal realizations”, defined as the (time-invariant) CTMCs in which all transition rates $R(s,s^{\prime})$ coincide with $m(s,s^{\prime})$ resp. $M(s,s^{\prime})$. These lumpabilities can be computed efficiently and are characterized both via value functions and satisfaction equivalence of formulas in a variant of the continuous stochastic logic CSL [9, 13]. However, there does not seem to be a direct connection to our notion of $(\varepsilon,\delta)$-bisimulation.
Another related line of research is perturbation theory for Markov chains. Given a CTMC $\mathcal{M}$ and a (slightly) perturbed $\mathcal{M}^{\prime}$, the goal is to bound the difference of (some) performance measures of the models. See, e.g., [43, 44, 45, 51, 58] or the overview article [52]. The bounds presented in the literature are usually on the total error between the stationary distributions of $\mathcal{M}$ and $\mathcal{M}^{\prime}$. Furthermore, they are oftentimes obtained under the assumption of drift conditions or rely on the ergodicity of the chain. This is in contrast to our work, which focuses on the componentwise difference in transient reachability probabilities up to time $t$.
@@ 2 Preliminaries
A continuous-time Markov chain (CTMC) is a tuple $\mathcal{M}=(S,P,E,s_{init},L)$ such that $\mathcal{D}_{\mathcal{M}}=(S,P,s_{init},L)$ is a DTMC, called the embedded or underlying DTMC of $\mathcal{M}$, and $E\colon S\to\mathbb{R}_{>0}$ is an exit rate function. We use the same notations for CTMCs as for DTMCs, and we let $\mathcal{M},\mathcal{N}$ range over CTMCs. Note that we do not exclude the possibility of $P(s,s)>0$, i.e., we allow self-loops in CTMCs. The residence time in a state $s$ of $\mathcal{M}$ is negative exponentially distributed with rate $E(s)$, so the probability to take any outgoing transition of $s$ until time $t\geq 0$ (including self-loops) is $1-e^{-E(s)\cdot t}$, and the probability to take a transition from $s$ to some specific state $s^{\prime}$ until $t$ is $P(s,s^{\prime})\cdot(1-e^{-E(s)\cdot t})$. $\mathbf{Q}\in\mathbb{R}^{|S|\times|S|}$ denotes the (infinitesimal) generator of $\mathcal{M}$. The entries of $\mathbf{Q}$ are given as $\mathbf{Q}_{i,j}=P(i,j)\cdot E(i)$ if $i\neq j$ and $\mathbf{Q}_{i,i}=-\sum_{j\neq i}~\mathbf{Q}_{i,j}$. For more information on CTMCs and DTMCs see, e.g., [14, 36, 39].
@@ 3.2 Definition of $(\varepsilon,\delta)$-Bisimulation
Let $\varepsilon<\frac{1}{2}$ and $\delta>0$. In Figure 2, $\sim_{\varepsilon,\delta}$ is the reflexive and symmetric closure of $\{(s_{0},s_{2}),(s_{0},s_{3}),(s_{1},s_{4}),(s_{2},s_{3})\}$. States $s_{1}$ and $s_{2}$ are not $(\varepsilon,\delta)$-bisimilar as they violate the $\delta$-condition: $|\ln(E(s_{1}))-\ln(E(s_{2}))|=2\delta$. Moreover, $s_{0}\nsim_{\varepsilon,\delta}s_{1}$ since, although the pair of states satisfies the $\delta$-condition, the $\varepsilon$-condition is violated: $P(s_{0},\{s_{2}\})=\frac{1}{2}>\varepsilon=P(s_{1},{\sim_{\varepsilon,\delta}}(\{s_{2}\}))+\varepsilon.$
Note that the steady state distributions of $(\varepsilon,\delta)$-bisimilar states can differ significantly: starting from $s_{0}$ the probability to reach $g$ (and stay there forever) is $2/3$, from $s_{2}$ it is $1$ and from $s_{3}$ it is 00, even though $s_{0},s_{2}$, and $s_{3}$ are pairwise $(\varepsilon,\delta)$-bisimilar.
@@ 3.3 Fundamental Properties of $\sim_{\varepsilon,\delta}$
By slightly adjusting a procedure proposed in [25] that constructs $\sim_{\varepsilon}$ for DTMCs by solving flow networks á la [10, 57], $\sim_{\varepsilon,\delta}$ can be computed in time polynomial in the number of states of $\mathcal{M}$.
@@
Furthermore, $(\varepsilon,\delta)$-bisimilarity in $\mathcal{M}$ implies $\tau$-bisimilarity (for a $\tau\geq 0$ that depends on $\varepsilon$ and $\delta$) in the uniformization $\mathrm{unif}(\mathcal{M},q)$ of $\mathcal{M}$ for $q\geq\max_{s\in S}E(s)$.
@@
For any given $\varepsilon\in(0,1),\delta>0$ and $\tau>0$ there is a CTMC $\mathcal{M}=\mathcal{M}(\varepsilon,\delta,\tau)$ with states $s,s^{\prime}$ such that $\{s,s^{\prime}\}$ is a block of a $\tau$-quasi lumpability on $\mathcal{M}$ but the pair $(s,s^{\prime})$ does neither satisfy the $\varepsilon$- nor the $\delta$-condition.
@@
We finish the section by showing how to “split” $(\varepsilon,\delta)$-bisimilarity into $(\varepsilon,0)$- and $(0,\delta)$-bisimilarity. More precisely, given $\mathcal{M}\sim_{\varepsilon,\delta}\mathcal{N}$ we construct CTMCs $\mathcal{M}^{\prime},\mathcal{N}^{\prime}$ with the same graph structure and such that $\mathcal{M}\sim_{\varepsilon,0}\mathcal{M}^{\prime}\sim_{0,\delta}\mathcal{N}^{\prime}\sim\mathcal{N}$. This decomposition makes it possible to treat the two parameters individually and, together with the additivity of $\sim_{\varepsilon,\delta}$, allows us to extend results shown for $(\varepsilon,0)$- and $(0,\delta)$-bisimilar states or chains to $(\varepsilon,\delta)$-bisimilar ones.
@@ 4 Bounding Timed Reachability Probabilities
By considering the DTMC $\mathrm{unif}(\mathcal{M},q)$ and applying the bound of Theorem 3.1 [18], which is possible because of the preservation of approximate bisimilarity in the uniformization (see Lemma 2), we derive an upper bound on the absolute difference of timed reachability probabilities of $(\varepsilon,\delta)$-bisimilar states.
@@
A disadvantage of the bounds in Proposition 3 (and Corollary 2) is that they converge to $1$ exponentially fast for $t\to\infty$, while the maximal total difference in timed reachability probabilities is usually much smaller for $(\varepsilon,\delta)$-bisimilar states. Intuitively, the bounds converge to $1$ because the bound of Theorem 3.1 used in their derivation always assumes the maximal possible error. However it does not consider, e.g., that for growing $t$ some of the probability mass already reached the goal state and can thus not contribute to the total error anymore. This disadvantage is particularly observable for $(0,\delta)$-bisimilar states, as in this case the actual error converges to 00 for $t\to\infty$ (see also Figures 6 and 7 in Section 5). In Section 5, we derive explicit formulas and better bounds for the absolute difference in timed reachability probabilities of $(0,\delta)$-bisimilar states.
@@ 5 The Effect of Changing Rates Only
Next we show how to obtain, for given $t$ and $\delta$, a maximal $N$ (that depends on $t,\delta$) such that $\mathrm{Diff}_{t}(\mathcal{M})\leq\mathrm{Diff}_{t}(\mathcal{E}_{N})$. Using the explicit form of $\mathrm{Diff}_{t}(\mathcal{E}_{N})$ from Proposition 4, this yields an easy-to-compute upper bound for $\mathrm{Diff}_{t}(\mathcal{M})$.
@@
Next, we present a way to compute the values of $p_{n}$ based on a spectral decomposition of the transition probability matrix $\mathbf{P}$ of CTMC $\mathcal{M}$. We start with the case of a diagonalizable $\mathbf{P}$, and afterwards deal with arbitrary $\mathbf{P}$ by using the Jordan canonical form [50]. Moreover, we use the so obtained explicit formulas for $p_{n}$ to obtain upper bounds for $\mathrm{Diff}_{t}(\mathcal{M})$. In contrast to the results so far, these bounds converge to 00 for $t\to\infty$, just like the actual difference does.
The transition probability matrix $\mathbf{P}$ of $\mathcal{M}$ is diagonalizable if there is a diagonal matrix $\mathbf{D}$, with diagonal elements corresponding to the eigenvalues of $\mathbf{P}$ (repeated according to their multiplicities), and a regular matrix $\mathbf{S}$ such that $\mathbf{P}=\mathbf{S}\mathbf{D}\mathbf{S}^{-1}$ [50]. It is well-known that every eigenvalue $\lambda_{i}$ of a stochastic matrix satisfies $|\lambda_{i}|\leq 1$, and that $\lambda_{1}=1$ is an eigenvalue of any such matrix. As we assume that an absorbing state is reached almost surely, $1$ is the only eigenvalue of modulus $1$, and it has a multiplicity $a_{\mathbf{P}}$ of at most $2$ (as there are at most two absorbing states in $\mathcal{M}$). Hence, we can w.l.o.g. assume that for the $m$ distinct eigenvalues of $\mathbf{P}$ we have $\lambda_{1}=1>|\lambda_{2}|\geq\ldots\geq|\lambda_{m}|$, and that the diagonal of $\mathbf{D}$ is in descending order w.r.t. the absolute values of these eigenvalues. From now on we denote the second largest (absolute value wise) eigenvalue of $\mathbf{P}$ by $\lambda$.
@@
By Proposition 6, $p_{k+1}=\sum_{j=a_{\mathbf{P}}+1}^{n}\mathbf{S}_{1,j}\cdot\mathbf{S}^{-1}_{j,n}\cdot(\lambda_{j}-1)\cdot\lambda_{j}^{k}$, so by the triangle inequality $p_{k+1}~\leq\sum_{j=a_{\mathbf{P}}+1}^{n}|\mathbf{S}_{1,j}\cdot\mathbf{S}^{-1}_{j,n}\cdot(\lambda_{j}-1)|\cdot|\lambda_{j}|^{k}$. Because $|\lambda_{j}|\leq|\lambda|$ for all $j>a_{\mathbf{P}}$ we get $p_{k+1}\leq|\lambda|^{k}\cdot\sum_{j=a_{\mathbf{P}}+1}^{n}C=|\lambda|^{k}\cdot(n-a_{\mathbf{P}})\cdot C$ for $C=\max_{i=a_{\mathbf{P}}+1,\ldots,n}|\mathbf{S}_{1,i}\cdot\mathbf{S}^{-1}_{i,n}\cdot(\lambda_{i}-1)|$. The claim follows from Theorem 5.1. ∎
@@
Let $\mathcal{M}^{\prime}=e^{0.1}\cdot\mathcal{M}$. In particular, $\mathcal{M}\sim_{0,0.1}\mathcal{M}^{\prime}$. We analyze the bounds from Propositions 5 and 7 w.r.t. $\mathcal{M},\mathcal{M}^{\prime}$ and $f$ by considering three scenarios that are parameterized in $\tau$ and $\mu$ [^2] We omit the bound based on uniformization from Proposition 3 as we have already seen in Example 5 that it does not perform well for $(0,\delta)$-bisimilar models. In the first scenario, $\tau_{1}=\frac{3}{4}$ and $\mu_{1}=\frac{1}{4}$, i.e., the arrival rate of tasks is three times higher than their service rate. In the second scenario, arrival and service rates coincide ($\tau_{2}=\frac{1}{2}=\mu_{2}$), while in the third scenario the service rate is twice the arrival rate, i.e., $\tau_{3}=\frac{1}{3}$ and $\mu_{3}=\frac{2}{3}$. We denote the CTMC corresponding to scenario $i\in\{1,2,3\}$ by $\mathcal{M}_{i}$. Note that for each $i$ the transition probability matrix $\mathbf{P}_{i}$ of $\mathcal{M}_{i}$ is diagonalizable.
@@
We observe that in each scenario the actual error and the bounds of Proposition 7 converge to 00 for $t\to\infty$, while the bound of Proposition 5 converges to $1$ for increasing $t$. The convergence rate to 00 of the former bound can, however, be quite slow. In, e.g., the third scenario, the second largest eigenvalue of $\mathbf{P}_{3}$ has modulus $0.9778$, causing a slow convergence of $|\lambda|^{k}$ to 00. On the other hand, in the first scenario $\mathcal{M}_{1}$ enters $f$ quickly, which is reflected in a second largest eigenvalue of $0.7334$ whose powers converge to 00 fast.
While the values of the bounds from Proposition 7 are, in particular for small $t$, significantly larger than the actual error, it is promising that their shape as a function of $t$ closely resembles the actual difference. This scaling of the bounds by a large factor is caused by the constant $C$, which is obtained by using the triangle inequality and a maximum. Refining this constant in future work might lead to bounds following the actual error much more closely.
@@
The decompositions of $\mathbf{P}$ proved in Propositions 6 and 8 also hold when the Markov chain contains more than $2$ (reachable) absorbing states, i.e, if $a_{\mathbf{P}}>2$. In an absorbing Markov chain, the multiplicity (both geometric and algebraic) of the eigenvalue $1$ is the number of absorbing states, and thus each Jordan block for eigenvalue $1$ has a size of $1$, causing them to get canceled out in the computations of the explicit forms of $p_{N+1}$.
@@
In Theorem 5.2 we have to exclude the case $|\lambda|=0$, i.e., the case that $\mathbf{P}$ only has eigenvalues 00 and $1$. This happens, e.g., if $\mathbf{P}$ (and hence $\mathcal{M}$) is acyclic. In an acyclic CTMC the goal state $g$ is, however, reached after at most as many steps as the length of the longest path in the chain. Therefore, if $\mathcal{M}$ is acyclic, $\mathrm{Diff}_{t}(\mathcal{M})$ can easily be computed exactly, for example by using the identity in Theorem 5.1 and computing the relevant values of $p_{n}=\mathbf{P}^{n+1}-\mathbf{P}^{n}$, or by applying methods like the ACE algorithm of [47] that allow computing the transient distribution functions of acyclic CTMCs directly.
@@
6 Bounds for Reward-Bounded Reachability Probabilities